OpenStudy (anonymous):

Help!! f(x+8) = (x + 8)2-5x(x+8)+1

myininaya (myininaya):

I'm having trouble figuring out what you need here.

myininaya (myininaya):

What are the instructions or directions?

OpenStudy (anonymous):

to solve it

OpenStudy (anonymous):

it ended up being x2 + 11x +25 but i dont know how

myininaya (myininaya):

? What are we solving for?

myininaya (myininaya):

Do you mean you want to simplify f(x+8)?

OpenStudy (anonymous):

yes thank you that helps

myininaya (myininaya):

\[(x+8)^2-5x(x+8)+1=(x^2+16x+64)-5x(x+8)+1\] \[=(x^2+16x+64)-5x^2-40x+1=x^2-5x^2+16x-40x+64+1\] \[-4x^2-24x+65\]

myininaya (myininaya):

without that extra x in there you would get what you got above

OpenStudy (anonymous):

what you want to find?

OpenStudy (anonymous):

I guess, people mostly ask us to find critical points for this type of questions. So, by solving RHS, you get \[−4x2−24x+65\] 1. Let us differentiate LHS and RHS. \[d/dx (f(x+8) ) = d/dx ( −4x2−24x+65) \] at max and min points, LHS = 0. so, equate RHS to 0 to find the points at which LHS=0; so the critical points. Note: I think, there must be some range given, so take only the critical points which are in that range. If range isn't given.. take them all. 2. Other critical points are where f(x+8) is undefined. as its a polynomial.. it is defined for all R. So, you should be getting 2 critical points.

OpenStudy (anonymous):

Well, if you want to find \[f(x) \] from it.. its very simple.. substitute so, the above equation \[f(x+8) = (x + 8)^2-5x(x+8)+1\] becomes \[f(x-8+8) = (x-8 + 8)^2-5(x-8)(x-8+8)+1\] \[=> f(x) = x^2 -5(x-8)(x)+1\] \[=> f(x) = x^2 -5x^2 +40x+1\] \[=> f(x) =-4x^2 +40(x)+1\] * please do check if there are any arithmetic errors (minor +/-).. However, the basic structure and method is right

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

I forgot to mention a word in my above answer [in 5th lin3] . "substitute x = x-8